OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 108, -18, 3, -1/2, ...

OpenStudy (amistre64):

is this an arithmetic or geometric sequence?

OpenStudy (amistre64):

in a geometric sequence, as long as the common ratio is less than 1 it should converge to zero.

OpenStudy (amistre64):

should prolly correct that as:\[|r|<1\]

OpenStudy (anonymous):

I don't really know... but the limit was 648/7.. which I don't understand

OpenStudy (amistre64):

from the looks of it 3r = 1/2 if r = 1/6 108/6 = 18 18/6 = 3 3/6 = 1/2 (1/2)/6 = 1/12 1/12 /6 = 1/72 as the terms continue, the value tends towards 0

OpenStudy (amistre64):

the limit of the sequence is therefore 0

OpenStudy (anonymous):

If the limits 0 why was the answer 648/7? or is it the same thing?

OpenStudy (amistre64):

im not sure what your 'answer' is unless your not looking at the right place for the answer maybe? if you can upload a screenshot or pic of the actual question itll help avoid misinterpretations.

OpenStudy (anonymous):

It may be an error made by my teacher - I'll consult with him. Thank you

OpenStudy (amistre64):

good luck